# WorldAPIs: The World Is Worth How Many APIs? A Thought Experiment

### v1

@@ Abstract
AI systems make decisions in physical environments through primitive actions or affordances that are accessed via API calls. While deploying AI agents in the real world involves numerous high-level actions, existing embodied simulators offer a limited set of domain-salient APIs. This naturally brings up the questions: how many primitive actions (APIs) are needed for a versatile embodied agent, and what should they look like?
We explore this via a thought experiment: assuming that wikiHow tutorials cover a wide variety of human-written tasks, what is the space of APIs needed to cover these instructions? We propose a framework to iteratively induce new APIs by grounding wikiHow instruction to situated agent policies. Inspired by recent successes in large language models (LLMs) for embodied planning, we propose a few-shot prompting to steer GPT-4 to generate Pythonic programs as agent policies and bootstrap a universe of APIs by 1) reusing a seed set of APIs; and then 2) fabricate new API calls when necessary. The focus of this thought experiment is on defining these APIs rather than their executability.
@@
NLRSE Workshop @ ACL’24
WorldAPIs: The World Is Worth How Many APIs?
A Thought Experiment
Jiefu Ou,  Arda Uzunoğlu,  Benjamin Van Durme,  Daniel Khashabi Johns Hopkins University {jou6, auzunog1, vandurme, danielk}@jhu.edu
@@ 1 Introduction
Figure 1: Top: WorldAPIs, the proposed thought experiment that takes a top-down approach. Starting from daily tasks with sequences of instruction steps in wikiHow, and a seed action space (API pool), we iteratively prompt LLMs to generate agent programs and add the induced (hallucinated) APIs in generated programs to the API pool.
Bottom: in contrast, most of the prior work in building embodied environments often adopts a bottom-up approach. The simulation and collection of instructions and programs are all based on a close set of predefined actions.
@@
Developing versatile and capable AI agents that follow natural language instructions for task execution in the real world has been an important research objective. A series of efforts (Ahn et al., 2022; Singh et al., 2022a; Song et al., 2022) have been made towards this goal, with a notable upsurge of work on harnessing the power of large language models (LLMs) to generate task plans and policies for embodied agents.
@@
The backbones of such success are benchmarks and simulations of embodied environments. (Anderson et al., 2018; Thomason et al., 2019a; Puig et al., 2018a; Shridhar et al., 2020; Ku et al., 2020). However, most of these existing embodied environments suffer from the limited hand-crafted action spaces. Such restraint diminishes action diversity, leading to the oversaturation of a small group of basic actions. For example, as shown in Figure 1, existing embodied simulators/environments often define a small and closed set of actions an agent can perform. As a result, the possible instructions it can carry out are highly constrained by the limited action space. However, carrying out tasks in real-world scenarios across different domains requires a much larger and more diverse action space. To bridge this gap between embodied agent learning and real-world deployment, it is natural to ask the following questions: To build a versatile embodied agent that can carry out daily tasks in the physical world, how many primitive actions (APIs) should such an agent be equipped with, and what do they look like?
@@
In this work, we make the first step towards answering these questions by proposing WorldAPIs3), a thought experiment that aims to approximate a realistic pool of agent-executable primitive actions (APIs) [^1] In this work we refer to APIs as interfaces to affordances of agents that could be deployed in the real world. Upon calling an API, the agent can carry out a corresponding physical primitive action. We use APIs and primitive actions and APIs interchangeably henceforth, the composition of which enables agents to accomplish diverse and realistic goals specified by high-level procedural natural language instructions. As shown in Figure 1, in contrast to prior work, we propose to induce an open action space from diverse and realistic procedural instructions.
@@
We conduct preliminary experiments (§5) to induce agent policies and action spaces on wikiHow [^2] https://www.wikihow.com, a rich semi-structured online resource of human-written procedural instructions covering daily tasks from a wide spectrum of domains. The set of APIs saturated after running the iterative induction pipeline on less than $0.5\%$ of wikiHow tutorials, yielding a self-contained pool of over 300 frequently-evoked APIs (§6.2). This provides an approximation of the lower bound of the primitive action space. Human and automatic evaluations (§6.1) demonstrate the proposed pipeline’s effectiveness in inducing APIs necessary to carry out diverse wikiHow instructions. Further comparison with existing embodied environments (Shridhar et al., 2020; Puig et al., 2018b) highlights the limited coverage of our induced APIs in existing environments (§6.2), thus encouraging further efforts to enrich the action space of embodied simulations.
@@ 2 Related Work
The literature deals with enabling agents to navigate and manipulate their environments. Although LLMs are shown to be effective foundations for plan generation (Song et al., 2023; Huang et al., 2022, 2023), embodied planning usually requires more than textual data due to its interactive nature. Thus, creating a reflection of the real world (e.g., simulation or benchmarks) has been an important aspect of the progress. These simulators enable interactive learning and evaluation on a proxy of the world.
Various simulators have been developed for simulating human activities (Kolve et al., 2017; Savva et al., 2019; Xia et al., 2018; Gan et al., 2021a; Nasiriany et al., 2024), each with distinct advantages. Yet, they are intrinsically confined to a set of interactions based on factors such as the feasibility of rendering actions, computational limitations, constraints of the physics engine, etc. These benchmarks for embodied reasoning (summarized in Table 1) can be split into two groups: (i) grounded in physical world and (ii) grounded in software world.
@@
ALFRED (Shridhar et al., 2020) 13 Household Tasks
VirtualHome (Puig et al., 2018b) 12 Household Tasks
TEACh (Padmakumar et al., 2021) 16 Household Tasks
LACMA (Yang et al., 2023) 10 Household Tasks
BEHAVIOR (Srivastava et al., 2021) 6 Household Tasks
BEHAVIOR-1K (Li et al., 2024) 6 Household Tasks
Mini-BEHAVIOR (Jin et al., 2023) 15 Household Tasks
House3D (Wu et al., 2018) 12 Navigation
R2R (Anderson et al., 2018) 6 Navigation
CVDN (Thomason et al., 2019b) 6 Navigation
HM3D (Ramakrishnan et al., 2021) 4 Navigation
ThreeDWorld (Gan et al., 2021b) 6 Navigation,Manipulation
CHAI (Misra et al., 2019) 5 Navigation,Manipulation
CALVIN (Mees et al., 2021) 12 Manipulation
Overcooked (Carroll et al., 2019) 6 Cooking
VRKitchen (Gao et al., 2019) 6 Cooking
KITchen (Younes and Asfour, 2024) 12 Household & Cooking
SmartPlay (Wu et al., 2023) 37 Games
@@
ToolAlpaca (Tang et al., 2023) 400 Function Call
API-Bank (Li et al., 2023) 53 Function Call
API-Bench (Patil et al., 2023) 1.6K Function Call
API-Pack (Guo et al., 2024) 11.2K Function Call
ToolBench (Xu et al., 2023) 232 Function Call
ToolLLM (Qin et al., 2023) 16.4K Function Call
@@
There is a broad range of works that focus on procedures, including reasoning on procedures (Zhang et al., 2020; Yang et al., 2021; Uzunoglu et al., 2024), generating procedures (Sakaguchi et al., 2021; Lyu et al., 2021), and so forth. Most of this literature is confined to unimodal tasks, and the multimodal approaches are limited (Wu et al., 2022; Zhou et al., 2023; Zellers et al., 2021). Similarly, the focus has been on individual tasks, even though collective benchmarks that cover different tasks exist (Uzunoglu and Şahin, 2023; Onoe et al., 2021; Wu et al., 2022). Therefore, most works are confined to shallow textual and visual representations. This leads to the lack of procedural data grounded in real-world tasks and scenarios. Similar to what we aspire to achieve, Puig et al. (2018a) utilizes wikiHow to convert high-level procedures to executable actions in virtual environments. However, its breadth is highly narrow due to its limited domain and predefined action space of 12 actions.
@@ 3.1 Collecting instructions from wikiHow
We leverage wikiHow, a prominent web platform with 200K+ professionally curated “how-to" tutorials across a diverse set of domains. Each wikiHow tutorial presents a sequence of instructional paragraphs in natural language that aim to accomplish a certain goal (see Figure 1 for an example). We follow prior work (Zhang et al., 2020; Zhou et al., 2022) to use the tutorial title as the goal, the paragraph headline as instruction steps, and the paragraph body as additional descriptions.
@@ 3.3 Principles for Translating Language Instructions to Pythonic Policies and Actions
Inspired by recent work (Liang et al., 2022; Singh et al., 2022a), we jointly induce primitive actions (APIs) and policies (Pythonic programs) via prompting LLMs with few-shot demonstrations (details in §4). The demonstrations provide information about the hypothetical environment to the LLMs: available objects, primitive actions (APIs), as well as how to interact with objects through API calls and state checking. We create an annotation guideline that defines the semantic formalism of objects and APIs for the hypothetical environment. We annotate the programs for a small set of wikiHow tutorials as seed demonstrations.
@@
The example in 1 shows a demonstration that translates the first wikiHow tutorial in the top-left of Figure 1, i.e. “How to Melt Chocolate in Microwave”. After specifying the tutorial (TASK) and numbered instruction steps (INSTRUCTIONS), the following PROGRAM lists the objects and primitive actions through import, specifies object relations with receptacles, and incorporates instruction steps and sub-steps in comments for policy specification.
TASK:
How to Melt Chocolate in Microwave
INSTRUCTIONS:
@@
PROGRAM:
# primitive APIs
from utils import find, grab, put, put_back ...
# objects
from objects_pool import chocolate_0, knife_0 ...
# object-object relation specification
faucet_0.receptacles.append(sink_0)
# program
def robot_program():
# 1. Chop the chocolate into small pieces with a serrated knife.
# find and grab chocolate
find(obj=chocolate_0)
grab(obj=chocolate_0)
# find and grab knife
find(obj=knife_0)
grab(obj=knife_0)
# find cutting board
find(obj=cutting_board_0)
# chop chocolate until it’s into small pieces
while chocolate_0.material_properties[‘form’] != ‘small pieces’:
chop(obj=chocolate_0, tool=knife_0, on=cutting_board_0)
put_back(obj=knife_0)
# 2. Place the chocolate into a microwave-safe bowl.
# execute the program
robot_program()
Listing 1: Our in-context demonstrations for decomposing wikiHow tasks into API calls.
@@
Actionable properties define the set of permitted APIs that the agent can evoke to interact with a particular object. For example, a grab API can take an object as the argument if and only if the object has an actionable property grabbable.
Material properties define the properties of an object that are not directly associated with any interaction but can be changed/affected as a result of participating in API calls. For example, the chocolate object in 1 with a material property {‘form’: ‘bar’} can be changed to {‘form’: ‘small pieces’} after participating in a chop API.
Object-object relationships can be specified via receptacles and receptacles_of, when necessary. We refer readers to §A.1 for the complete definition of WorldObject.
@@
While there are numerous potential ways to define the action spaces, we seek to strike a balance in our definition among 1) stylistically and semantically similar to existing embodied environments (Shridhar et al., 2020); 2) simple and informative such that LLMs can effectively leverage its parametric knowledge of code generation to predict agent policies as programs for unseen instructions; 3) easily extendable where LLMs are capable of hallucinating new APIs that are consistent to the available APIs specified in the demonstrations. Specifically, we focus on the following aspects:
Base APIs: To form a comparable basis with existing work, a set of base APIs that are commonly evoked in carrying out wikiHow instructions are provided, such as find, put, open_obj, turn_on, etc. We provide details of base APIs in §A.2.1.
Granularity: It is challenging to precisely define the right abstraction level of actions. Nevertheless, in our definitions, we aim to strike a balance between the following two principles: (Further details on these principles in §A.2.2)
  • Avoiding overly-abstract actions: We prefer to have tasks broken into actions that directly interact with the objects. For example, given the task of “dry your clothes with a drier,” we avoid distilling it to a single dry action. Instead, break up this task into sub-steps (e.g., “the dryer can be opened and turned on to perform drying”), each appropriate action.
  • Avoiding too low-level actions: We prefer to avoid actions that involve low-level physics. For instance in 1, given the instruction substep “chop chocolate until it ’s into small pieces”, we will define chop(obj=chocolate_0, tool=knife_0, on=cutting_board_0), instead of move_held_object(moveMagnitude=0.1).
State checking and feedback loop: Similar to prior work (Singh et al., 2022b), in our demonstrations we implement steps where the agent collects feedback from the environments by checking the object state (material properties) and acting accordingly. (details in §A.2.3)

### v2

@@ Abstract
AI systems make decisions in physical environments through primitive actions or affordances that are accessed via API calls. While deploying AI agents in the real world involves numerous high-level actions, existing embodied simulators offer a limited set of domain-salient APIs. This naturally brings up the questions: how many primitive actions (APIs) are needed for a versatile embodied agent, and how should they look like?
We explore this via a thought experiment: assuming that wikiHow tutorials cover a wide variety of human-written tasks, what is the space of APIs needed to cover these instructions? We propose a framework to iteratively induce new APIs by grounding wikiHow instruction to situated agent policies. Inspired by recent successes in large language models (LLMs) for embodied planning, we propose a few-shot prompting to steer GPT-4 to generate Pythonic programs as agent policies and bootstrap a universe of APIs by 1) reusing a seed set of APIs; and then 2) fabricate new API calls when necessary. The focus of this thought experiment is on defining these APIs rather than their excitability.
@@ 1 Introduction
Figure 1: Top: WorldAPIs, the proposed thought experiment that takes a top-down approach. Starting from daily tasks with sequences of instruction steps in wikiHow, and a seed action space (API pool), we iteratively prompt LLMs to generate agent programs and add the induced (hallucinated) APIs in generated programs to the API pool. Bottom: in contrast, most of the prior work in building embodied environments often adopts a bottom-up approach. The simulation and collection of instructions and programs are all based on a close set of predefined actions.
@@
Developing versatile and capable AI agents that follow natural language instructions for task execution in the real world has been an important research objective. A series of efforts (Ahn et al. 2022; Singh et al. 2022; Song et al. 2023) have been made towards this goal, with a notable upsurge of work on harnessing the power of large language models (LLMs) to generate task plans and policies for embodied agents.
@@
The backbones of such success are benchmarks and simulations of embodied environments. (Anderson et al. 2018; Thomason, Gordon, and Bisk 2019; Puig et al. 2018; Shridhar et al. 2020; Ku et al. 2020). However, most of these existing embodied environments suffer from the limited hand-crafted action spaces. Such restraint diminishes action diversity, leading to the oversaturation of a small group of basic actions. For example, as shown in Figure 1, existing embodied simulators/environments often define a small and closed set of actions an agent can perform. As a result, the possible instructions it can carry out are highly constrained by the limited action space. However, carrying out tasks in real-world scenarios across different domains requires a much larger and more diverse action space. To bridge this gap between embodied agent learning and real-world deployment, it is natural to ask the following questions: To build a versatile embodied agent that can carry out daily tasks in the physical world, how many primitive actions (APIs) should such an agent be equipped with, and what do they look like?
@@
In this work, we make the first step towards answering these questions by proposing WorldAPIs3), a thought experiment that aims to approximate a realistic pool of agent-executable primitive actions (APIs) [^1] In this work we refer to APIs as interfaces to affordances of agents that could be deployed in the real world. Upon calling an API, the agent can carry out a corresponding physical primitive action. We use APIs and primitive actions and APIs interchangeably henceforth, the composition of which enables agents to accomplish diverse and realistic goals specified by high-level procedural natural language instructions. The thought experiment assumes an embodied agent to be deployed in a hypothetical world consisting of objects and primitive action space with predefined properties. The agent is tasked to perform a wide variety of daily tasks specified through natural language instructions, in the form of wikiHow tutorials. To complete these tasks, the agent needs to evoke and compose primitive actions (APIs) from the action space to interact with objects. As shown in Figure 1, in contrast to prior work, we propose to induce an open action space from diverse and realistic wikiHow procedural instructions.
@@
We conduct preliminary experiments (§5) to induce agent policies and action spaces on wikiHow  [^2] https://www.wikihow.com, a rich semi-structured online resource of human-written procedural instructions covering daily tasks from a wide spectrum of domains. The set of APIs saturated after running the iterative induction pipeline on less than $0.5\%$ of wikiHow tutorials, yielding a self-contained pool of over 300 frequently-evoked APIs (§6.2). This provides an approximation of the lower bound of the primitive action space. Human and automatic evaluations (§6.1) demonstrate the proposed pipeline’s effectiveness in inducing APIs necessary to carry out diverse wikiHow instructions. Further comparison with existing embodied environments (Shridhar et al. 2020; Puig et al. 2018) highlights the limited coverage of our induced APIs in existing environments (§6.2), thus encouraging further efforts to enrich the action space of embodied simulations.
@@ 2 Related Work
This literature focuses on enabling agents to navigate and manipulate their environments. While LLMs are shown to be effective at planning (Song et al. 2023; Huang et al. 2022, 2023), embodied planning usually requires more than textual data due to its interactive nature. Thus, creating a reflection of the real world (e.g., simulation or benchmarks) has been an important aspect of the progress. These simulators enable interactive learning and evaluation on a proxy of the world.
Various simulators have been developed for simulating human activities (Kolve et al. 2017; Xia et al. 2018; Nasiriany et al. 2024), each with distinct advantages. Yet, they are intrinsically confined to a set of interactions based on factors such as the feasibility of rendering actions, computational limitations, constraints of the physics engine, etc. These benchmarks for embodied reasoning (summarized in Table 1) can be split into two groups: (i) grounded in physical world and (ii) grounded in software world.
@@
ALFRED (Shridhar et al. 2020) 13 Household Tasks
VirtualHome (Puig et al. 2018) 12 Household Tasks
TEACh (Padmakumar et al. 2021) 16 Household Tasks
LACMA (Yang et al. 2023) 10 Household Tasks
BEHAVIOR (Srivastava et al. 2021) 6 Household Tasks
BEHAVIOR-1K (Li et al. 2024) 6 Household Tasks
Mini-BEHAVIOR (Jin et al. 2023) 15 Household Tasks
House3D (Wu et al. 2018) 12 Navigation
R2R (Anderson et al. 2018) 6 Navigation
CVDN (Thomason et al. 2019) 6 Navigation
HM3D (Ramakrishnan et al. 2021) 4 Navigation
ThreeDWorld (Gan et al. 2021) 6 Navigation,Manipulation
CHAI (Misra et al. 2019) 5 Navigation,Manipulation
CALVIN (Mees et al. 2021) 12 Manipulation
Overcooked (Carroll et al. 2019) 6 Cooking
VRKitchen (Gao et al. 2019) 6 Cooking
KITchen (Younes and Asfour 2024) 12 Household & Cooking
SmartPlay (Wu et al. 2023) 37 Games
@@
ToolAlpaca (Tang et al. 2023) 400 Function Call
API-Bank (Li et al. 2023) 53 Function Call
API-Bench (Patil et al. 2023) 1.6K Function Call
API-Pack (Guo et al. 2024) 11.2K Function Call
ToolBench (Xu et al. 2023) 232 Function Call
ToolLLM (Qin et al. 2023) 16.4K Function Call
@@
There is a broad range of works that focus on procedures, including reasoning on procedures  (Zhang, Lyu, and Callison-Burch 2020; Uzunoglu, Safa, and Şahin 2024), generating procedures (Sakaguchi et al. 2021; Lyu, Zhang, and Callison-Burch 2021), and so forth. Most of this literature is confined to unimodal tasks, and the multimodal approaches are limited (Zhou et al. 2023; Zellers et al. 2021). Similarly, the focus has been on individual tasks, even though collective benchmarks that cover different tasks exist (Uzunoglu and Şahin 2023; Onoe et al. 2021) Therefore, most works are confined to shallow textual and visual representations. This leads to the lack of procedural data grounded in real-world tasks and scenarios. Similar to what we aspire to achieve, Puig et al. (2018) utilizes wikiHow to convert high-level procedures to executable actions in virtual environments. However, its breadth is highly narrow due to its limited domain and predefined action space of 12 actions.
@@ 3.1 Collecting instructions from wikiHow
We leverage wikiHow, a prominent web platform with 200K+ professionally curated “how-to” tutorials across a diverse set of domains. Each wikiHow tutorial presents a sequence of instructional paragraphs in natural language that aim to accomplish a certain goal (see Figure 1 for an example). We follow prior work (Zhang, Lyu, and Callison-Burch 2020; Zhou et al. 2022) to use the tutorial title as the goal, the paragraph headline as instruction steps, and the paragraph body as additional descriptions.
@@ 3.3 Principles for Translating Language Instructions to Pythonic Policies and Actions
Inspired by recent work (Liang et al. 2022; Singh et al. 2022), we jointly induce primitive actions (APIs) and policies (Pythonic programs) via prompting LLMs with few-shot demonstrations (details in §4). The demonstrations provide information about the hypothetical environment to the LLMs: available objects, primitive actions (APIs), as well as how to interact with objects through API calls and state checking. We create an annotation guideline that defines the semantic formalism of objects and APIs for the hypothetical environment. We annotate the programs for a small set of wikiHow tutorials as seed demonstrations.
@@
The example in Listing 1 shows a demonstration that translates the first wikiHow tutorial in the top-left of Figure 1, i.e. “How to Melt Chocolate in Microwave”. After specifying the tutorial (TASK) and numbered instruction steps (INSTRUCTIONS), the following PROGRAM lists the objects and primitive actions through import, specifies object relations with receptacles, and incorporates instruction steps and sub-steps in comments for policy specification.
Listing 1 Our in-context demonstrations for decomposing wikiHow tasks into API calls.
1TASK:
2How to Melt Chocolate in Microwave
3INSTRUCTIONS:
@@
6...
7PROGRAM:
8# primitive APIs
9from utils import find, grab, put, put_back ...
10# objects
11from objects_pool import chocolate_0, knife_0 ...
12# object-object relation specification
13faucet_0.receptacles.append(sink_0)
14...
15# program
16def robot_program():
17 # 1. Chop the chocolate into small pieces with a serrated knife.
18 # find and grab chocolate
19 find(obj=chocolate_0)
20 grab(obj=chocolate_0)
21 # find and grab knife
22 find(obj=knife_0)
23 grab(obj=knife_0)
24 # find cutting board
25 find(obj=cutting_board_0)
26 # chop chocolate until it’s into small pieces
27 while chocolate_0.material_properties[‘form’] != ‘small pieces’:
28 chop(obj=chocolate_0, tool=knife_0, on=cutting_board_0)
29 put_back(obj=knife_0)
30 # 2. Place the chocolate into a microwave-safe bowl.
31 ...
32# execute the program
33robot_program()
@@
Actionable properties specify the APIs an agent can invoke to interact with an object. For example, the grab API can only target objects with the actionable property grabbable.
Material properties refer to the inherent attributes of objects that are not directly tied to interactions but can be altered through API calls. For instance, the chocolate object in Listing1 with the material property {‘form’: ‘bar’} may change to {‘form’: ‘small pieces’} after a chop API call.
Object-object relationships can be specified via receptacles and receptacles_of, when necessary. We refer readers to apendix on our arXiv draft for the complete definition of WorldObject. We refer readers to §A.1 for the complete definition of WorldObject.
@@
While there are numerous potential ways to define the action spaces, we seek to strike a balance in our definition among 1) stylistically and semantically similar to existing embodied environments (Shridhar et al. 2020); 2) simple and informative such that LLMs can effectively leverage its parametric knowledge of code generation to predict agent policies as programs for unseen instructions; 3) easily extendable where LLMs are capable of hallucinating new APIs that are consistent to the available APIs specified in the demonstrations. Specifically, we focus on the following aspects:
Base APIs: To ensure comparability with prior work, we provide a set of commonly used APIs for executing wikiHow instructions, such as find, put, open_obj, and turn_on. Details are in §A.2.
Granularity: Defining the right abstraction level for actions is challenging. Nevertheless, in our definitions, we aim to strike a balance between the following two principles: (Further details on these principles in §A.2)
  • Avoiding overly-abstract actions: We prefer to have tasks broken into actions that directly interact with the objects. For example, given the task of “dry your clothes with a drier,” one straightforward approach is distilling it to a single dry action. However, we argue that it is not a desired approach to create the action space since it creates APIs at a problematic level of granularity from an action perspective - drying something with a dryer requires a totally different set of actions to drying something with iron. Instead, break up this task into sub-steps (e.g., “the dryer can be opened and turned on to perform drying”), each appropriate action.
  • Avoiding too low-level actions: We prefer to avoid actions that involve low-level physics. As this degrades the action space into just a handful of control APIs with complicated spatial-temporal argument realization, which is infeasible in our hypothetical environment without explicit spatial relation specification, and harms the reusability of learned APIs in practice. For instance in Listing 1, given the instruction substep “chop chocolate until it ’s into small pieces”, we will define chop(obj=chocolate_0, tool=knife_0, on=cutting_board_0), instead of move_held_object(moveMagnitude=0.1).
State checking and feedback loop: Similar to prior work (Singh et al. 2022), in our demonstrations we implement steps where the agent collects feedback from the environments by checking the object state (material properties) and acting accordingly (details in §A.2)